# Social Choice for AI Alignment: Dealing with Diverse Human Feedback

### v1

@@ Abstract
Foundation models such as GPT-4 are fine-tuned to avoid unsafe or otherwise problematic behavior, so that, for example, they refuse to comply with requests for help with committing crimes or with producing racist text. One approach to fine-tuning, called reinforcement learning from human feedback, learns from humans’ expressed preferences over multiple outputs. Another approach is constitutional AI, in which the input from humans is a list of high-level principles. But how do we deal with potentially diverging input from humans? How can we aggregate the input into consistent data about “collective” preferences or otherwise use it to make collective choices about model behavior? In this paper, we argue that the field of social choice is well positioned to address these questions, and we discuss ways forward for this agenda, drawing on discussions in a recent workshop on Social Choice for AI Ethics and Safety held in Berkeley, CA, USA in December 2023.
@@ 1 Introduction
Over the past year, reinforcement learning from human feedback (RLHF) has played a key role in making large language models (LLMs) more capable and controllable (Christiano et al., 2017; Ziegler et al., 2019). RLHF is now the primary strategy that leading AI companies such as OpenAI (OpenAI, 2023), Anthropic (Anthropic, 2023), Meta (Meta, 2023), and Google (Google, 2023) use to align pretrained LLM models with human values. However, RLHF faces many limitations and concrete challenges (Casper et al., 2023; Lambert & Calandra, 2023), including unrepresentative data (Prabhakaran et al., 2021; Feffer et al., 2023), unrealistic models of human decision-making (Hong et al., 2022; Freedman et al., 2021; Siththaranjan et al., 2023; Lambert et al., 2023), and insufficient modeling of human diversity (Kirk et al., 2023; Freedman et al., 2023). We hold the position that core ideas from social choice theory (Arrow, 2012; Fishburn, 1973; Kelly, 1988; Brandt et al., 2015)—primarily concerning whose preferences should be integrated into decisions and how this should be done—are needed to solve many of the open problems facing RLHF.
@@
While models that are solely pretrained on internet data may produce repetitive or harmful text, RLHF enables training models to follow instructions (Ouyang et al., 2022) and produce helpful and “harmless” outputs (Bai et al., 2022a) based on human judgments. RLHF gathers example outputs from an LLM that has been pretrained to predict a text corpus. Next, humans are asked to select the outputs that best meet specified criteria (such as being “helpful” or “unbiased”). Humans may also manually write the outputs to be compared, but due to cost, human input is often limited to these comparative judgements. These judgments, often called preferences, are then used to fine-tune the LLM to produce more desirable outputs. From a social choice perspective, this method raises several critical questions: Which humans are asked to judge models? What criteria do they use? How are their judgments combined? And how do their expressed judgments relate to their actual preferences?
@@
Constitutional AI (CAI), which involves reinforcement learning from AI feedback (RLAIF), is an alternate approach that directly addresses some of these questions (Bai et al., 2022b). Humans produce a “constitution” that explicitly specifies principles to guide the LLM training process. The LLM is then trained to align with this constitution. However, we must still decide who has input on the constitution and how it is constructed. Bai et al. (2022b) construct their constitution “in a fairly ad-hoc way […] for research purposes”, but developing safe and ethical AI requires a more principled approach, as exemplified in Ganguli et al. (2023) or announced in OpenAI (2024). How then should one aggregate diverse preferences into a representative constitution?
@@
Social choice theory has long studied similar questions, and by taking into account its lessons, one can avoid making naïve mistakes and reinventing the wheel. In this paper, we argue that tools and theories from social choice should be applied to these open problems, in particular in RLHF, to help bridge challenging design problems to sociotechnical questions (Dobbe et al., 2021). Specifically, we demonstrate how such tools can be used to begin addressing which humans should provide input or feedback, what type of feedback they should provide, and how that feedback should be aggregated and used. We also highlight areas in which new work is required to extend social choice to new problems unique to training safe and ethical AI.
@@
There are a number of advantages to addressing these problems in a principled way. First, it is likely to result in a fairer system that takes into account the input or feedback of a broader group of people. Second, there are reasons to believe that this will result in generally more accurate feedback about questions of truthfulness. Indeed, there is a significant body of literature on “epistemic democracy”—voting to settle questions about facts (Pivato, 2017). Intuitively, having input from a more diverse group of people makes it less likely that something important is missed. Third, it will likely result in broader buy-in into the system. For example, important issues such as political biases of LLMs (Motoki et al., 2023) have been hypothesized to emerge from the finetuning phase that follows pretraining (Rozado, 2024).
@@
One may also have concerns about this approach; for example, is feedback from a diverse group of people going to be inconsistent and consequently result in inconsistent behavior from the system? Social choice theory provides a number of examples where naïve aggregation of preferences or judgments leads to choices in the aggregate that seem irrational, such as cyclical preferences (Schwartz, 2018) or logically inconsistent conclusions (List & Pettit, 2002). Then again, social choice theory also provides the tools for thinking about such issues and preventing them.
@@
The field of social choice is not new to computer scientists; computational social choice (Brandt et al., 2015) is by now a well-studied topic, with a dedicated biennial workshop since 2006. However, while many of the researchers in this area affiliate with the AI community, there has not yet been much work connecting computational social choice to the alignment of modern AI systems.
@@
In the remainder of this paper, we first give background on value alignment, RLHF, and social choice. Then we discuss a number of questions at the intersection of these topics. We believe that significant further research is required to answer each of these questions well and that good answers to them are needed to build AI systems in a responsible way based on potentially diverging feedback from multiple stakeholders. In contrast, ad-hoc approaches to these questions may result in systems that fail to represent their stakeholders well, that fail to address important issues, that marginalize significant groups of stakeholders, and that create a basis for conflict between groups of people or the multiple AI systems that represent them.
@@ 2 Background
Our proposed research agenda requires background on topics that have so far been studied by mostly disjoint communities. A reader familiar with some of these topics can skip the corresponding subsections.
@@ 2.1 Value Alignment
As advanced AI systems become increasingly capable, it becomes critical that they act in a way that aligns with human and societal values (Gabriel, 2020). There are many approaches to value alignment, including theoretical work to define formal games that AI agents must align with humans to solve (Shah et al., 2020), empirical investigation of the relationship between neural network activations and morally relevant output features (Zou et al., 2023), and evaluations of the ethical behavior of state-of-the-art models (Pan et al., 2023). RLHF is a particularly popular approach to value alignment, but it faces many limitations in its current form.
@@ 2.2 Reinforcement Learning from Human Feedback
Preference data collection
The first step in RLHF is to generate and evaluate a dataset of model outputs $\mathcal{Y}$. In vanilla RLHF, humans are then shown paired completions $\{y_{0},y_{1}\}\in\mathcal{Y}\times\mathcal{Y}$ to prompts $x\in\mathcal{X}$ of these outputs and asked to select which output $p\in\{y_{0},y_{1}\}$ they prefer from each pair (Christiano et al., 2017; Lee et al., 2021). Other RLHF variants require humans to rank or provide scores for groups of outputs (Ziegler et al., 2019; Ouyang et al., 2022), and many additional variations exist (Wu et al., 2023).
Reward model training
The next step is to fit a parameterized reward model $\varrho_{\theta}\>:\>\mathcal{Y}\rightarrow\mathbb{R}$. For LLMs, the reward model is typically a neural network with weights $\theta$. RLHF methods assume that there is a ground-truth reward function $\varrho_{\theta^{*}}$ that the human preferences reflect up to probabilistic noise. The reward model is then optimized to match the likelihoods of the human preferences observed in the data. If the training data comes from diverse sources, this implicitly amounts to a rather intransparent form of preference aggregation (Siththaranjan et al., 2023).
Optimizing the policy with RL
The final step is to use reinforcement learning to train a policy that maximizes rewards from the reward model. This involves many design decisions—which RL algorithm to use, how to regularize the updates, and whether to gather further online feedback during training. See Uc-Cetina et al. (2023) for a survey of methods and limitations for using RL to train LLMs.
@@ 2.3 Alternate Preference-based Fine-tuning Objectives
Due to the instability of popular RL optimizers such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), several novel techniques for optimizing a model based on a collected preference dataset have been proposed. Rafailov et al. (2023) introduce Direct Preference Optimization (DPO), which recasts RLHF to converge to the optimum reward modeling solution by directly optimizing a loss on the preference label dataset, rather than sampling online from the LLM policy or training an explicit reward model. Another variant emerged to remove the dependency on pairwise data. Ethayarajh et al. (2023) propose a loss function termed Kahneman-Tversky Optimization (KTO) that enables learning a policy from unpaired preferences. The authors further claim that the effectiveness of various losses for RLHF depends on the properties they share with proposed human utility functions (Kahneman & Tversky, 1979).
@@ 2.4 Constitutional AI
Bai et al. (2022b) further explore the design space by introducing Constitutional AI (CAI), which relies on RL from AI Feedback (RLAIF). RLAIF is a set of techniques for using an AI model to augment or generate feedback data in the form of pairwise preferences or other signals (Lee et al., 2023; Sharma et al., 2024; Castricato et al., 2024). By employing a human-written set of principles, which they term a constitution, they use a separate LLM to generate artificial preference and instruction data that can be used for model fine-tuning. A constitution $\mathcal{C}$ is made up of a set of written principles $c_{i}$ that indicate specific aspects to focus on during a critique phase. The instruction data, which is largely out of the scope of this paper, is curated by repeatedly sampling a principle $c_{i}$ and asking the model to revise the current completion $y_{k}^{0}$ to the prompt $x_{k}$. This yields a series of instruction variants $\{y_{k}^{0},y_{k}^{1},\cdots,y_{k}^{n}\}$ from the principles $\{c_{i_{0}}^{0},c_{i_{1}}^{1},\cdots,c_{i_{n-1}}^{n-1}\}$ used for critique at each step. The final data point is the prompt $x_{k}$ with the final completion $y_{k}^{n}$, for some suitable $n$.
The preference data is constructed in a similar, yet simpler way by using a subset of principles from the constitution $\mathcal{C}$ as context for a feedback model. The feedback model is presented with a prompt $x$, a set of principles $\{c_{0},\cdots,c_{n}\}$, and two completions $y_{0}$ and $y_{1}$ labeled as answers (A) and (B) from a previous RLHF dataset. The feedback models’ probability of outputting either (A) or (B) is recorded as a training sample for the reward model, as discussed in Section 2.2.
@@ 2.5 Social Choice
Modern social choice theory began in the 1950s with Arrow’s Impossibility Theorem (Arrow, 1951). [^1] For its long prehistory, see McLean & Urken (1995). Arrow considered the problem of aggregating multiple individuals’ preferences—in the form of complete and transitive rankings of some set of alternatives—into a social preference, subject to a list of normative desiderata. In particular, Arrow assumed that the aggregation function should be defined for any family of individual preferences to be aggregated (Universal Domain); that the outputted social preference relation should be complete and transitive, like individual preferences, in which case the aggregation function is called a social welfare function; that the social preference between two alternatives $A$ and $B$ should depend only on individual preferences between $A$ and $B$ (Independence of Irrelevant Alternatives); and that unanimous individual preference for $A$ over $B$ should imply social preference for $A$ over $B$ (Pareto). Arrow proved that if there are at least three alternatives, then the only aggregation functions satisfying these desiderata are dictatorships: there is one individual $d$ such that no matter what others prefer, if $d$ strictly prefers $A$ to $B$, then the social preference ranks $A$ over $B$ as well. [^2] Mishra (2023) applies Arrow’s Theorem to RLHF. A similar theorem (see Taylor 2005, § 1.3) holds for social choice functions where, instead of asking for a social ranking of alternatives, we more modestly ask for just a set of choice-worthy alternatives. This also includes the special case of social choice functions that always pick a single winner.
Arrow’s Theorem stimulated a huge literature exploring the consequences of weakening Arrow’s desiderata (see, e.g., Campbell & Kelly 2002, Holliday & Pacuit 2020, and references therein). The general takeaway is that for ordinal preference aggregation, in order to avoid dictatorships and related pathologies such as oligarchies and vetoers, one must weaken the Independence of Irrelevant Alternatives (IIA) and allow the social preference between two alternatives to depend in part on individual preferences involving other alternatives. With this freedom to relax IIA comes a vast proliferation of alternative methods of aggregating individual preferences (see, e.g., Brams & Fishburn 2002; Zwicker 2016; Pacuit 2019 and the voting methods implemented in the Preferential Voting Tools library). Figure 1 gives an example in which three well-known methods disagree. The costs and benefits of these and other methods are systematically studied from different angles (axiomatic, computational, empirical, etc.) in social choice theory.
@@
Figure 1: Individual rankings on the left (4 voters submit the ranking $ABC$, 4 submit $ACB$, etc.) lead to different aggregated rankings on the right, depending on the aggregation rule. Borda Count gives an alternative 0 points for each voter who ranks it last, 1 point for each voter who ranks it second, and 2 points for each voter who ranks it first; alternatives are then ordered by descending score. Instant Runoff ranks $C$ last since $C$ has the fewest first-place rankings; then, after removing $C$ from all voters’ rankings, $B$ has the fewest first-place rankings, so $B$ is in second and $A$ is in first. For Ranked Pairs, notice there is a majority cycle: a majority of voters prefer $A$ to $B$, a majority prefer $B$ to $C$, and a majority prefer $C$ to $A$; but the smallest majority margin of victory is for $A$ over $B$, so we reverse this majority preference, yielding $BCA$.
Since Arrow, social choice theory has grown to study aggregation not only of individuals’ preferences, both ordinal and cardinal (d’Aspremont & Gevers, 2002), but also of their approvals of alternatives (Laslier & Sanver, 2010), grades given to alternatives (Balinski & Laraki, 2010), judgments about propositions (Grossi & Pigozzi, 2022), subjective probabilities for propositions (Dietrich & List, 2016), and other types of objects (Rubinstein & Fishburn, 1986). In the following, we discuss some of the aggregation problems that might arise in the context of AI alignment.

### v2

# Social Choice Should Guide AI Alignment in Dealing with Diverse Human Feedback

@@ Abstract
Foundation models such as GPT-4 are fine-tuned to avoid unsafe or otherwise problematic behavior, such as helping to commit crimes or producing racist text. One approach to fine-tuning, called reinforcement learning from human feedback, learns from humans’ expressed preferences over multiple outputs. Another approach is constitutional AI, in which the input from humans is a list of high-level principles. But how do we deal with potentially diverging input from humans? How can we aggregate the input into consistent data about “collective” preferences or otherwise use it to make collective choices about model behavior? In this paper, we argue that the field of social choice is well positioned to address these questions, and we discuss ways forward for this agenda, drawing on discussions in a recent workshop on Social Choice for AI Ethics and Safety held in Berkeley, CA, USA in December 2023.
@@ 1 Introduction
Recently, reinforcement learning from human feedback (RLHF) has become the primary strategy that leading AI companies such as OpenAI (OpenAI, 2023), Anthropic (Anthropic, 2023), Meta (Meta, 2023), and Google (Google, 2023) use to make pretrained large language models (LLMs) more capable and controllable (Christiano et al., 2017; Ziegler et al., 2019) and to align them with human values. However, RLHF faces many limitations and challenges (Casper et al., 2023; Lambert & Calandra, 2023), including unrepresentative data (Prabhakaran et al., 2021; Feffer et al., 2023), unrealistic models of human decision-making (Hong et al., 2022; Freedman et al., 2021; Siththaranjan et al., 2023; Lambert et al., 2023), and insufficient modeling of human diversity (Kirk et al., 2023; Freedman et al., 2023) which may lead to political bias (Motoki et al., 2023; Rozado, 2024). We propose that ideas from social choice theory (Arrow, 2012; Fishburn, 1973; Kelly, 1988; Brandt et al., 2015)—e.g., concerning whose preferences should be integrated into decisions and how this should be done—are needed to solve many of these open problems.
@@
While models that are solely pretrained on internet data may produce repetitive or harmful text, RLHF trains models to follow instructions (Ouyang et al., 2022) and produce helpful and “harmless” outputs (Bai et al., 2022a) based on human judgments. RLHF gathers example outputs from a pretrained LLM or examples written by humans. Next, humans are asked to select the outputs that best meet specified criteria (such as being “helpful” or “unbiased”). These judgments, often called preferences, are then used to fine-tune the LLM. From a social choice perspective, this method raises several critical questions: Which humans are asked to judge outputs? What criteria do they use? How are their judgments combined? And how do their expressed judgments relate to their actual preferences?
@@
Constitutional AI (CAI), which involves reinforcement learning from AI feedback (RLAIF), is an alternative addressing some of these questions (Bai et al., 2022b). Humans produce a “constitution” that specifies principles that the LLM is trained to align with. However, one must still decide who has input on the constitution and how it is constructed. Bai et al. (2022b) construct it “in a fairly ad-hoc way […] for research purposes”, but developing safe and ethical AI requires a more principled approach as in Ganguli et al. (2023) or OpenAI (2024). How then should one aggregate diverse preferences into a representative constitution?
@@
Social choice theory has long studied similar questions. This position paper argues that methods from social choice should be applied to address questions such as which humans should provide input, what type of feedback should be collected, and how it should be aggregated and used. By taking into account the lessons from social choice theory, one can avoid naïve mistakes and reinventing the wheel, while leveraging feedback to address challenging design problems (Dobbe et al., 2021). We also highlight areas in which new work is required to extend social choice to new problems unique to training safe and ethical AI.
@@
There are several advantages to addressing the above problems in a principled way. First, it will likely result in a fairer system, taking into account the input of a broader group of people. Second, it promises to give generally more accurate feedback about questions of truthfulness, as suggested by a significant body of literature on “epistemic democracy”—voting to settle questions about facts (Pivato, 2017). Intuitively, having input from diverse people makes it less likely that something important is missed. Third, it will likely result in broader buy-in into the system.
@@
One may have concerns about this approach; for example, will feedback from diverse people be inconsistent and result in inconsistent system behavior? Social choice theory provides examples where naïve aggregation of preferences or judgments leads to seemingly irrational collective choices, such as cyclical preferences (Schwartz, 2018) or logically inconsistent conclusions (List & Pettit, 2002). Then again, social choice theory also provides the tools for thinking about such issues and preventing them.
@@
Social choice is not new to computer scientists; computational social choice (Brandt et al., 2015) is a well-studied topic, with a biennial workshop since 2006. However, while many of those researchers affiliate with the AI community, there has not yet been much work connecting computational social choice to the alignment of modern AI systems.
@@
In the following, we first give background on value alignment, RLHF, and social choice. Then we discuss several questions at their intersection. We think that significant further research is needed to answer these questions well and that good answers are needed in order to build AI systems in a responsible way based on potentially diverging feedback from multiple stakeholders. In contrast, ad-hoc approaches may result in systems that fail to represent their stakeholders well, that marginalize significant groups of stakeholders, and that create a basis for conflict between groups of people or the multiple AI systems that represent them.
@@ 2 Background
Our proposed research agenda requires background from so far mostly disjoint communities. Readers familiar with some of the following can skip the corresponding subsections.
@@ 2.1 Value Alignment
As AI systems become more capable, it becomes critical that they act in alignment with human and societal values (Gabriel, 2020). Many approaches to value alignment exist, such as, e.g., formal games in which AI agents must align with humans to solve them (Shah et al., 2020), empirics on the relation between neural network activations and morally relevant output features (Zou et al., 2023), and evaluations of the ethical behavior of LLMs (Pan et al., 2023). RLHF is a particularly popular but so far limited approach.
@@ 2.2 Reinforcement Learning from Human Feedback
RLHF begins with generating and evaluating a dataset of model outputs $\mathcal{Y}$. In vanilla RLHF, humans are then shown paired completions $\{y_{0},y_{1}\}\in\mathcal{Y}\times\mathcal{Y}$ to prompts $x\in\mathcal{X}$ and asked which output $y\in\{y_{0},y_{1}\}$ they prefer (Christiano et al., 2017; Lee et al., 2021). Other RLHF variants ask humans to rank or provide scores for groups of outputs (Ziegler et al., 2019; Ouyang et al., 2022), and many additional variations exist (Wu et al., 2023).
The next step is to fit a parameterized reward model $\varrho_{\theta}\colon\mathcal{Y}\rightarrow\mathbb{R}$. For LLMs, $\varrho_{\theta}$ is typically a neural network with weights $\theta$. RLHF methods assume that there is a ground-truth reward function $\varrho_{\theta^{*}}$ that the human preferences reflect (up to noise). The reward model is then optimized to match the likelihoods of the human preferences observed in the data. If the training data comes from diverse sources, this implicitly amounts to a rather intransparent or flawed form of preference aggregation (Siththaranjan et al., 2023; Xu et al., 2023; Chakraborty et al., 2024; Ge et al., 2024).
The final step is to use reinforcement learning to train a policy that maximizes rewards from the reward model. This involves many design decisions—which RL algorithm to use, how to regularize the updates, and whether to gather further online feedback during training. See Uc-Cetina et al. (2023) for a survey on the use of RL to train LLMs.
@@ 2.3 Alternate Preference-Based Fine-Tuning Objectives
Since common RL methods like PPO (Schulman et al., 2017) can be unstable, novel techniques for optimizing a model based on collected preference data have been proposed. Rafailov et al. (2023) introduce Direct Preference Optimization (DPO), which recasts RLHF to converge to the best solution by directly optimizing a loss on the preference label dataset, rather than sampling online from the LLM policy or training an explicit reward model. Another variant emerged to remove the dependency on pairwise data. Ethayarajh et al. (2023) propose a loss function termed Kahneman–Tversky Optimization (KTO) that enables learning a policy from unpaired preferences. The authors further claim that the effectiveness of various losses for RLHF depends on the properties they share with proposed human utility functions (Kahneman & Tversky, 1979).
@@ 2.4 Constitutional AI
Bai et al. (2022b) further explore the design space with Constitutional AI (CAI), which relies on RL from AI Feedback (RLAIF). RLAIF is a larger set of techniques for using AI to augment or generate feedback data, including pairwise preferences (Lee et al., 2023; Sharma et al., 2024; Castricato et al., 2024). By employing a human-written set of principles, which they term a constitution, Bai et al. (2022b) use a separate LLM to generate artificial preference and instruction data used for fine-tuning. A constitution $\mathcal{C}$ is a set of written principles indicating specific aspects to focus on during a critique phase. The instruction data is curated by repeatedly sampling a principle $c_{i}\in\mathcal{C}$ and asking the model to revise its latest output $y^{i}$ to the prompt $x$ to align with $c_{i}$. This yields a series of instruction variants $\{y^{0},y^{1},\cdots,y^{n}\}$ from the principles $\{c_{0},c_{1},\cdots,c_{n-1}\}$ used for critique. The final data point is the prompt $x$ together with the final completion $y^{n}$, for some $n$.
The preference data is constructed in a similar, yet simpler way by using a subset of principles from $\mathcal{C}$ as context for a feedback model. The feedback model is presented with a prompt $x$, a set of principles $\{c_{0},\cdots,c_{n}\}$, and two completions $y_{0}$ and $y_{1}$ labeled as answers (A) and (B) from a previous RLHF dataset. The feedback models’ probability of outputting either (A) or (B) is recorded as a training sample for the reward model, as discussed in Section 2.2.
@@ 2.5 Social Choice
Modern social choice theory began in the 1950s with Arrow’s Impossibility Theorem (Arrow, 1951; McLean & Urken, 1995). Arrow considered the problem of aggregating multiple individuals’ preferences—in his case rankings—into a social preference, subject to some desiderata. In particular, he required the aggregation function to be defined for any family of individual preferences; that the social preference relation be complete and transitive (it is then called a social welfare function); that the social preference between alternatives $A$ and $B$ should depend only on individual preferences between $A$ and $B$ (Independence of Irrelevant Alternatives, IIA); and that unanimous individual preference for $A$ over $B$ should imply social preference for $A$ over $B$. Arrow proved that the only aggregation functions satisfying these desiderata are dictatorships: there is one individual $d$ such that no matter what others prefer, if $d$ prefers $A$ to $B$, then the social preference ranks $A$ above $B$ as well. [^1] Mishra (2023) applies Arrow’s Theorem to RLHF. A similar theorem (see Taylor 2005, § 1.3) holds for social choice functions where, instead of asking for a social ranking, we ask for just a single winner, or even just a set of choice-worthy alternatives.
Arrow’s Theorem stimulated a huge literature exploring the consequences of weakening his desiderata (see, e.g., Campbell & Kelly 2002, Holliday & Pacuit 2020). The general takeaway is that for ordinal preference aggregation, in order to avoid dictatorships, oligarchies and vetoers, one must weaken IIA and allow the social preference between $A$ and $B$ to depend in part on preferences involving other alternatives. This allows for many alternative methods of aggregating individual preferences (see, e.g., Brams & Fishburn 2002; Zwicker 2016; Pacuit 2019 and the voting methods in the Preferential Voting Tools library). Figure 1 gives an example in which three well-known methods disagree. The costs and benefits of these and other methods are systematically studied from various angles (axiomatic, computational, empirical, etc.) in social choice theory.
@@
Figure 1: Individual rankings on the left (4 voters say $ABC$, 4 say $ACB$, etc.) lead to different aggregations on the right, depending on the aggregation rule. Borda Count gives an alternative 0, 1, or 2 points for each voter who ranks it last, second, or first, respectively; alternatives are then ordered by score. Instant Runoff ranks $C$ last since $C$ has the fewest first-place rankings; after removing $C$, $B$ has the fewest first-place rankings, so $B$ is in second and $A$ first. For Ranked Pairs, notice there is a majority cycle: a majority of voters prefer $A$ to $B$, a majority prefer $B$ to $C$, and a majority prefer $C$ to $A$; the smallest margin of victory is for $A$ over $B$, so we drop this majority preference, yielding $BCA$.
Since Arrow, social choice theory has grown to study aggregation not only of individuals’ preferences, both ordinal and cardinal (d’Aspremont & Gevers, 2002), but also of their approvals of alternatives (Laslier & Sanver, 2010), grades given to alternatives (Balinski & Laraki, 2010), judgments about propositions (Grossi & Pigozzi, 2022), subjective probabilities for propositions (Dietrich & List, 2016), and other types of objects (Rubinstein & Fishburn, 1986). Recent work has also started to explore social choice methods for RLHF (Song et al., 2024; Swamy et al., 2024; Dai & Fleisig, 2024; Ge et al., 2024). In the following, we discuss some of the aggregation problems that might arise in the context of AI alignment.